What is the solution to solve for x and y using elimination method in the double equation 2x-y=1; x+3y=4. and.
2x-4y=26; 5x+3y=26

Answer by Alan3354(69443)

solve for x and y using elimination method in the double equation 2x-y=1
x+3y=4 times 2 --> 2x + 6y = 8
---
2x - y = 1
2x +6y = 8
------------------ Subtract
-7y = -7
y = 1
---
2x - y = 1
2x - 1 = 1
2x = 2
x = 1
==========================
2x-4y=26
5x+3y=26
I would multiply the 1st eqn by 3 and the 2nd by 4, then add them.
================
PS It's not a "double equation" it's 2 equations.